Your firm needs a computerized machine tool lathe which costs $96,818 and requires $11,726 in maintenance expense for each year of its 5-year life. After five years, this machine will be replaced. The machine falls into the MACRS (Links to an external site.) 5-year class life category. Assume a tax rate of 25 percent and a discount rate of 12 percent. If the lathe can be sold for $14,799 at the end of year 5, what is the after-tax salvage value? Below are the MACRS rates for the 5-year recovery period. DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ENTER YOUR ANSWER TO THE NEAREST DOLLAR (e.g. 1250).

Year MACRS Rate

1 20.00%

2 32.00%

3 19.20%

4 11.52%

5 11.52%

6 5.76%
Answer #1

Written down value of asset after Year 5 = 96,818*5.76% = $5,576.7168

Selling Price = $14,799

Gain on Sale = $9,222.2832

Tax on gain = 9,222.2832*25% = $2,305.5708

After tax salvage value = Selling price – Tax

= 14,799 – 2305.5708

= $12,493.4292

i.e. $12,493